OpenStudy (anonymous):

Can some one help me with this problem? i will post it once i get a response

OpenStudy (tester97):

what level of math do you need help with?

OpenStudy (anonymous):

10th

OpenStudy (tester97):

alg 2?

OpenStudy (anonymous):

no geometry

OpenStudy (tester97):

oh sorry i havent did geometry in an while maybe i can try to help whats the question?

OpenStudy (anonymous):

it'll take time to type, one second

OpenStudy (anonymous):

ok so im just having trouble with getting my answer i know all the steps, or maybe im doing a step wrong. the question is Find values of x and for which ABCD must be a parallelogram. 5y - 8 = y + 24 5y+y= 6y 8+24= 32 6/32 = 5.333333 (I know something is wrong with the calculations)

OpenStudy (ajprincess):